Given the vectors v = -2i + 5j and w = 3i + 4j, determine
#1 1/2 v
#2 w - v
#3 length of w
#4 the unit vector for v

Answer by Alan3354(21583)

Given the vectors v = -2i + 5j and w = 3i + 4j, determine
#1 1/2 v
= -i + (5/2)j
-------------
#2 w - v
= 5i - j
----------
#3 length of w
= sqrt(3^2 + 4^2) = 5
----------------
#4 the unit vector for v
= v/length of v
= v/sqrt(2^2 + 5^2) = v/sqrt(29) = vsqrt(29)/29
= - 2i*sqrt(29)/29 + 5j*sqrt(29)/29